# Scale-Dropout: Estimating Uncertainty in Deep Neural Networks Using Stochastic Scale

### v1

@@ V-B Designing Spintronic-Based Scale Dropout Module
Figure 2: SOT-Based Scale-Dropout Module.
@@ V-C Proposed Spintronics-based CIM Architecture
Recent state-of-the-art (SOTA) CNN topologies, e.g., ResNet and DenseNet, use skip connections. In CIM, skip-connection can be implemented by selectively routing the output signals through the crossbars and summing them with digital circuits. Since layer-by-layer computations are sequential, signals for these connections can be stored in a buffer memory until the computation of the following layers is completed.
@@ VI-A2 OOD Dataset
We used six additional OOD datasets to evaluate the efficacy of our method in dealing with data uncertainty. : 1. Gaussian noise ($\hat{\mathcal{D}}_{1}$): Each pixel of the image is generated by sampling random noise from a unit Gaussian distribution, $\mathbf{x}\sim\mathcal{N}(0,1)$, 2. Uniform noise ($\hat{\mathcal{D}}_{2}$): Each pixel of the image is generated by sampling random noise from a uniform distribution, $\mathbf{x}\sim\mathcal{U}(0,1)$, 3. CIFAR-10 with Gaussian noise ($\hat{\mathcal{D}}_{3}$): Each pixel of the CIFAR-10 images is corrupted with Gaussian noise, 4. CIFAR-10 with uniform noise ($\hat{\mathcal{D}}_{4}$): Each pixel of the CIFAR-10 images is corrupted with uniform noise, 5. SVHN: Google street view house numbers dataset [42], and 6. STL10: a dataset containing images from the popular ImageNet dataset [43]. Each of these OOD datasets contains $8000$ images, and the images have the same dimensions as the original CIFAR-10 dataset ($32\times 32$ pixels). During the evaluation phase,
@@ VI-B Evaluated Topologies and Training setting
All weights and activations of models for classification tasks are binarized (1-bit model). The activations of biomedical semantic segmentation models are quantized to 4 bits, but their weights are kept binary. As stated previously, semantic segmentation tasks are more difficult and, therefore, require slightly more bit precision at the activation for accurate predictions. We have used the activation quantization algorithm proposed in [48] to quantize the activation to 4 bits. Since 1-bit weights are still maintained, no modifications to the crossbar structure are necessary. Only peripheral modifications, such as an ADC with an increased bit resolution, are required.
@@ VI-C Evaluation Metrics
(Epistemic) Uncertainty estimation of the models is evaluated on predictive variations, entropy, and confidence interval with $K=95\%$ based on Equations 19,21, respectively. Out-of-distribution data is detected as:
@@ VI-D Architectural Simulation
To carry out the architectural simulation, we first obtained the circuit specifications for the peripheral blocks, as outlined in Section V. We then simulated each component of the architecture independently to gauge its energy use. Both the crossbar array and the Scale-Dropout module were analyzed using SPICE to assess their energy consumption. The use of high-resistance SOT devices [49], in conjunction with the binary nature of the network, serves to reduce the overhead related to the peripheral elements. The Accumulator-Adder, Comparator, and Averaging circuits were synthesized using the Synopsys Design Compiler, leveraging the TSMC 40 nm low-power PDK-based standard IP libraries. For the CIM operation, decoding and sensing were assessed at the circuit-array level using NVsim [50]. To achieve this, we modified the NVsim simulator to accommodate multiple active cells, thus simulating CIM operation accurately. Additionally, we substituted the single-bit sense amplifiers with multi-bit ADCs. Performance metrics for each discrete component are shown in Table II.
@@
VI-E Predictive Performance
VI-E1 Comparison With State-of-the-Art Algorithms
The predictive performance of our method is comparable to SOTA binary Bayesian NN methods, as shown in Table VI on a range of CNN architectures, including VGG, ResNet-18 and ResNet-20, evaluated on the CIFAR-10 dataset. In the worst case, the predictive performance is $1.45\%$ below the SpinDrop [29] method for the VGG topology. Here, we assumed that there are no device variations in the spintronics-based scale dropout module. For a fair comparison, we have used the same network size as those used in their work. However, the hardware implementation for our solution may lead to a smaller area and a better power performance product owing to a simpler spintronics-based dropout module design. In our analysis, we have used layer-dependent adaptive dropout rates (See Section III-D) to scale dropout.
Furthermore, our proposed method achieves an improvement of up to $6.74\%$ in inference accuracy compared to the SOTA point estimate BNN algorithm. However, since our method is built on top of the IR-Net BNN algorithm [22], predictive performance should be comparable to their approach. As depicted in Table VI, the predictive performance is in the worst case $0.18\%$ lower, which is negligible. Similarly, accuracy is comparable to the full precision model, depicting that our method in general does not increase quantization error. Note that in the full precision model, the ReLU function is used as the activation for the convolution layers, while our proposed method uses the activation function $\operatorname{sign}(x)$ for all layers where activations are applied.
For biomedical image segmentation tasks, the proposed method outperforms the full-precision MC-Dropout method by up to $6.4$% in terms of IoU score. In the worst scenario, our method results in a $69.69$% reduction in the IoU score for the breast cancer dataset. Additionally, our approach outperforms the MC-Dropout method in most other metrics. Table IV presents a summary of the results.
VI-E2 Impact of MC Runs on Inference Accuracy
We observed that using Monte Carlo sampling ($T$ forward pass) for Bayesian inference generally enhances predictive performance across all datasets. For example, the inference accuracy of the ResNet-20 model increases from $84.63\%$ to $86.05\%$. In our evaluation, twenty samples (T = 20) for the larger model and fifty samples (T = 50) for the smaller model were used for Bayesian inference. Fig. 7 shows that the proposed method requires a smaller number of samples, with inference accuracy plateaus around $T=20$ to $50$. In comparison, MC-Dropout and MC-DropConnect methods require $100$, and $90$ Monte Carlo sampling, respectively, to achieve maximum inference accuracy [51]. Therefore, our method requires around $50$ less Monte Carlo sampling, which leads to 50×50$\times$50 × less XNOR and bit-counting operation, energy consumption, and latency for each Bayesian inference result. We assume the same NN topology, hardware architecture, and memory device technology for fair comparison.
VI-E3 Performance of BayBNN with Spin-ScaleDrop
TABLE III: Evaluation of the inference accuracy of the proposed Spintronics-based Scale-Dropout. The models trained with the variations in dropout probability $p$ increased from 0×0$\times$0 × to 3×3$\times$3 ×. Here, 0×0$\times$0 × variation represents the ideal scenario without variation.
Topologie Baseline (algorithmic) Trained w/ Spin-ScaleDrop Trained w/o Spin-ScaleDrop
Var. 0×0$\times$0 × Var. 1×1$\times$1 × Var. 2×2$\times$2 × Var. 3×3$\times$3 × Var. 0×0$\times$0 × Var. 1×1$\times$1 × Var. 2×2$\times$2 × Var. 3×3$\times$3 ×
ResNet-18 91.52% 91.77% 91.78% 91.77% 91.71% 91.61% 91.65% 91.59% 91.58%
VGG 90.45% 90.45% 90.52% 90.52% 90.55% 90.26% 90.28% 90.26% 90.30%
In the other case, the NN is trained considering the variation in the dropout module (see Section V-A). In this case, unlike in the previous case, there is a slight improvement in predictive performance (up to $+0.26\%$) for both models compared to the baseline. Variation in the dropout probability leads to more stochastically during Bayesian inference, and as a result, accuracy improves slightly. The results are summarized in Table III.
VI-F Uncertainty Estimation
Figure 5: Detecting Distribution Shift on CIFAR-10: a) A scatter and b) 95% confidence interval of 100 forward passes of the softmax input (logits) and output for Scale-Dropout VGG topology. Uniform noise of increasing strength is added to a randomly sampled image of a ship (levelled as 8). Although the model uncertainty is extremely high (best observed in colour), the input for images 5 through 12 is classified as either a truck (levelled as 9) or a bird (levelled as 3).
Figure 6: The effect of distribution shift of inference images on inference accuracy and predictive entropy (uncertainty estimate). Images are continuously rotated to introduce distribution shifts.
TABLE IV: The analysis of the proposed Scale-Dropout BayBNN on three biomedical segmentation datasets on SOTA topologies. A Dropout probability of 20% is used with 20 Monte Carlo samples for Bayesian inference.
U-Net Skin Cancer MC-Dropout 32/32 95.05% 81.67% 98.86% 89.91% 84.74% 98.67% 95.74%
Proposed 1/4 96.75% 88.07% 99.4% 93.66% 92.31% 98.31% 95.04%
Bayesian SegNet Breast Cancer MC-Dropout 32/32 97.83% 71.82% 96.40% 83.21% 78.23% 99.28% 88.86%
SpinDrop 1/4 97.47% 68.17% 95.48% 80.13% 74.25% 99.18% 87.02%
Proposed 1/4 97.69% 69.69% 96.69% 82.14% 77.19% 99.21% 87.76%
Bayesian SegNet COVID-19 Lung CT MC-Dropout 32/32 99.51% 72.88% 99.39% 84.32% 86.99% 99.7% 81.8%
Proposed 1/4 99.55% 74.43% 99.69% 85.34% 85.84% 99.76% 84.85%
Figure 8: The outcomes of semantic segmentation and uncertainty estimations for the Skin-Cancer, COVID-19 Lung-CT and Breast Cancer datasets. Each row comprises the input image followed by the ground truth, prediction mask, correctness mask, and uncertainty mask for both the MC-Dropout and proposed MC-Scale Dropout method. The correctness mask is a binary representation of accurate and inaccurate predictions. The uncertainty mask is the normalised [0,1] uncertainty mask derived from twenty Monte Carlo samples. For prediction masks, a 0.5 threshold is used. On the correctness and uncertainty masks, correct and certain regions are depicted in white. Similarly, a region that is incorrect or uncertain is displayed in black.
Detecting Distribution Shift
On the other hand, we have performed another experiment with all images of the CIFAR-10 dataset on the VGG model continuously rotated up to $90^{\circ}$. It can be seen in Fig.6 that as the images are rotated, the inference accuracy decreases and the predictive entropy increases from the starting entropy. Our method is compared against deterministic as well as common uncertainty estimation techniques, namely MC-Dropout [10] and Deep Ensemble [8] with five randomly initialized models. The trend of decrease in inference accuracy is similar for all models. However, our proposed MC Scale-Dropout method produces significantly more predictive entropy compared to other methods. Consequently, our method can produce better uncertainty estimates compared to related works even with a 1-bit model.
Note that in the MC-Dropout model, dropout is applied to the extracted features from the convolutional layers to achieve similar inference accuracy. If dropout is applied to all the layers, the predictive entropy increase but inference accuracy decrease significantly, e.g., by more than 3%. Our approach provides a good balance between uncertainty estimates without any degradation in accuracy.
TABLE V: Evaluation of the proposed MC-Scale Dropout method in detecting OOD across various topologies. All the models are trained on CIFAR-10 dataset.
VGG Proposed $\mathbf{95.36\%}$ $\mathbf{95.40\%}$ $\mathbf{96.19\%}$ $\mathbf{88.73\%}$ $\mathbf{98.02\%}$ $\mathbf{85.24\%}$
MC-Spatial Dropout [35] $99.99\%$ $100\%$ $92.9\%$ $78.91\%$ $99.81\%$ $100\%$
SpinBayes [31] $99.86\%$- $94.35\%$- $80.33\%$ $89.31\%$
ResNet-18 Proposed $\mathbf{100\%}$ $\mathbf{100\%}$ $\mathbf{97.49\%}$ $\mathbf{77.77\%}$ $\mathbf{91.53\%}$ $\mathbf{78.61\%}$
MC-Spatial Dropout [35] $100\%$ $100\%$ $100\%$ $92.26\%$ $99.98\%$ $97.39\%$
ResNet-20 Proposed $\mathbf{96.51\%}$ $\mathbf{100\%}$ $\mathbf{90.34\%}$ $\mathbf{93.55\%}$ $\mathbf{100\%}$ $\mathbf{99.8\%}$
MC-Spatial Dropout [35] $97.2\%$ $100\%$ $90.79\%$ $87.94\%$ $99.03\%$ $99.81\%$
Detecting Out-of-distribution Data
As depicted in Table V, our proposed method can achieve a detection rate of OOD of up to $100\%$ across various model architectures and six different OOD datasets ($\hat{\mathcal{D}}_{1}$ through $\hat{\mathcal{D}}_{6}$). There are some variations in OOD detection rates across different architectures for the same OOD dataset. However, even in these cases, our method can consistently achieve a high OOD detection rate, with the lowest detection rate being $77.77\%$ on the ResNet-18 model with $\hat{\mathcal{D}}_{4}$ dataset. However, when the threshold for SoftMax confidence increases from $95\%$ to $99\%$, the OOD detection rate in the dataset $\hat{\mathcal{D}}_{4}$ improved to $81.78\%$, an $\sim 4\%$ improvement. Compared to MC-Spatial Dropout and SpinBayes methods, the OOD detection rates are generally similar. In the worst case, the OOD detection rate is $\sim 14\%$ lower for the VGG topology on the $\hat{\mathcal{D}}_{6}$ dataset. Therefore, the results indicate that the proposed MC-Scale Dropout method is a robust and reliable solution to OOD detection across diverse model architectures and datasets.
Epistemic Uncertainty of Semantic Segmentaion
@@ VI-F Uncertainty Estimation
TABLE VI: Predictive Performance of the proposed MC-SparialDropout method in comparison with SOTA methods on CIFAR-10.
@@
Spatial-SpinDrop [35] 1/1 No $84.71\%$
Proposed 1/1 Yes $\mathbf{86.04\%}$
@@
SpinDrop [29] 1/1 Yes $91.95\%$
@@
Proposed 1/1 Yes $\mathbf{90.45\%}$
VI-G Hardware Overhead Analysis

### v2

@@ V-B Designing Spintronic-Based Scale Dropout Module
Figure 2: Spin Scale-Dropout Module based on SOT MTJ.
@@ V-C Proposed Spintronics-based CIM Architecture
Recent state-of-the-art CNN topologies, e.g., ResNet and DenseNet, use skip connections. In Computing in Memory architectures, skip-connection can be implemented by selectively routing the output signals through the crossbars and summing them with digital circuits. Since layer-by-layer computations are sequential, signals for these connections can be stored in a buffer memory until the computation of the following layers is completed.
@@ VI-A2 OOD Dataset
We used six additional OOD datasets to evaluate the efficacy of our method in dealing with data uncertainty. : 1. Gaussian noise ($\hat{\mathcal{D}}_{1}$): Each pixel of the image is generated by sampling random noise from a unit Gaussian distribution, $\mathbf{x}\sim\mathcal{N}(0,1)$, 2. Uniform noise ($\hat{\mathcal{D}}_{2}$): Each pixel of the image is generated by sampling random noise from a uniform distribution, $\mathbf{x}\sim\mathcal{U}(0,1)$, 3. CIFAR-10 with Gaussian noise ($\hat{\mathcal{D}}_{3}$): Each pixel of the CIFAR-10 images is corrupted with Gaussian noise, 4. CIFAR-10 with uniform noise ($\hat{\mathcal{D}}_{4}$): Each pixel of the CIFAR-10 images is corrupted with uniform noise, 5. SVHN: Google street view house numbers dataset [48], and 6. STL10: a dataset containing images from the popular ImageNet dataset [49]. Each of these OOD datasets contains $8000$ images, and the images have the same dimensions as the original CIFAR-10 dataset ($32\times 32$ pixels).
@@ VI-B Evaluated Topologies and Training setting
All weights and activations of models for classification tasks are binarized (1-bit model). The activations of biomedical semantic segmentation models are quantized to 4 bits, but their weights are kept binary. As stated previously, semantic segmentation tasks are more difficult and therefore, they require slightly more bit precision at the activation for accurate predictions. We have used the activation quantization algorithm proposed in [54] to quantize the activations to 4 bits. Since 1-bit weights are still maintained, the crossbar structure does not need to be modified. In fact, only peripheral modifications, such as ADC with increased bit resolution are required.
@@ VI-C Evaluation Metrics
(Epistemic) Uncertainty estimation of the models is evaluated on predictive variations, entropy, and confidence interval with $K=95\%$ based on Equations 21,23, respectively. Out-of-distribution data is detected as:
@@ VI-D Architectural Simulation
To carry out the architectural simulation, we first obtained the circuit specifications for the peripheral blocks, as outlined in Section V. We then independently simulated each component of the architecture to gauge its energy utilization. Both the crossbar array and the Spin Scale-Dropout module were analyzed using an electrical simulator such as the Simulation Program for Integrated Circuit (SPICE), to assess their energy consumption. The use of high-resistance SOT devices [55], in conjunction with the binary nature of the network, serves to reduce the overhead related to peripheral elements.
The Accumulator-Adder, Comparator, and Averaging circuits were synthesized using the Synopsys Design Compiler, leveraging the TSMC 40 nm low-power Process Design Kit (PDK). For the CIM operation, decoding and sensing were assessed at the circuit-array level using NVsim (NonVolatile memory simulator) [56]. To achieve this, we modified the NVsim simulator to accommodate multiple active cells, thus simulating CIM operation accurately. Additionally, we substituted the single-bit sense amplifiers with multi-bit ADCs. Performance metrics for each discrete component are shown in Table II.
@@
VII Evaluation
@@
TABLE III: Predictive performance of the proposed MC-SparialDropout method in comparison with SOTA methods on CIFAR-10. The accuracy closest to the MC-Dropout is in bold, and the number in the bracket shows the standard deviations of the accuracy after different repetitions.
@@
MC-Dropout(HTanh) [10] 32/32 Yes $86.94\%$
Spatial-SpinDrop [16] 1/1 Yes $84.71\%$
Proposed 1/1 Yes $\mathbf{86.04\%(\pm 0.039)}$
@@
MC-Dropout (HTanH) [10] 32/32 Yes $89.49\%$
MC-Dropout (ReLU) [10] 32/32 Yes $91.64\%$
MC-DropConnect [61] 32/32 Yes $91.36\%$
SpinDrop [17] 1/1 Yes $\mathbf{91.95\%}$
@@
Proposed 1/1 Yes ${90.45\%(\pm 0.052)}$
VII-A Predictive Performance
VII-A1 Comparison With State-of-the-Art Algorithms
The predictive performance of our method is comparable to the SOTA binary Bayesian NN methods, as shown in Table III on a range of CNN architectures, including VGG, ResNet-18, and ResNet-20, evaluated on the CIFAR-10 dataset. In the worst case, the predictive performance is $1.45\%$ below the SpinDrop [17] method for the VGG topology. Here, we assumed that there are no device variations in the spintronics-based scale Dropout module. For a fair comparison, we used the same network size as those used in their work. However, the hardware implementation of our solution may lead to a smaller area and a better power-performance product owing to a simpler spintronics-based Dropout module design. In our analysis, we have used layer-dependent adaptive Dropout rates (See Section III-D) to scale Dropout. The low variance in inference accuracy (numbers in brackets) shows the stability of the proposed approach.
In terms of the activation function, the proposed binary BayNN uses the Sign(.) function, which is an approximation of the hard Tanh function. In this case, the proposed method performs similarly to the MC-Dropout method. However, in the case of the ReLU activation function in the MC-Dropout model, the accuracy of the MC-Dropout model increases. Thus, the difference between the proposed method and the MC-Dropout increases to around $\sim 1\%$.
Furthermore, our proposed method achieves an improvement of up to $6.74\%$ in inference accuracy compared to the SOTA point estimate BNN algorithm. However, since our method is built on top of the IR-Net BNN algorithm [23], predictive performance should be comparable to their approach. As depicted in Table III, the predictive performance is in the worst case $0.18\%$ lower, which is negligible. Similarly, accuracy is comparable to the full precision model, depicting that our method in general does not increase quantization error. Note that in the full precision model, the ReLU function is used as the activation for the convolution layers, while our proposed method uses the activation function $\operatorname{sign}(x)$ for all layers where activations are applied.
For biomedical image segmentation tasks, the proposed method outperforms the full-precision MC-Dropout method by up to $6.4$% in terms of IoU score. In the worst scenario, our method results in a $69.69$% reduction in the IoU score for the breast cancer dataset. Additionally, our approach outperforms the MC-Dropout method in most other metrics. Table V presents a summary of the results.
VII-A2 Impact of MC Runs on Inference Accuracy
We observed that using Monte Carlo sampling ($T$ forward pass) for Bayesian inference generally enhances predictive performance across all datasets. For example, the inference accuracy of the ResNet-20 model increases from $84.63\%$ to $86.05\%$. In our evaluation, twenty samples (T = 20) for the larger model and fifty samples (T = 50) for the smaller model were used for Bayesian inference. Fig. 7 (a) shows that the proposed method requires a smaller number of samples, with inference accuracy plateaus around $T=20$ to $50$. In comparison, MC-Dropout and MC-DropConnect methods require $100$, and $90$ Monte Carlo sampling, respectively, to achieve the maximum inference accuracy, as reported in [61] for CIFAR-10. In our experiment (see Fig. 7 (b)), we observed that the MC-DropConnect method plateaus at $100$ Monte Carlo runs, and the MC-Dropout method plateaus at $200$ Monte Carlo runs on the same model (VGG) and dataset. Therefore, our method requires up to $180$ less Monte Carlo sampling, leading to 10×10$\times$10 × less XNOR and bit-counting operation, energy consumption, and latency for each Bayesian inference result. We assume the same NN topology, hardware architecture, and memory device technology for a fair comparison. However, it should be noted that $T$ at which accuracy plateaus can vary from task to task and from model to model.
VII-A3 Performance of BayBNN with Spin-ScaleDrop
TABLE IV: Evaluation of the inference accuracy of the proposed Spintronics-based Scale-Dropout with variations in the Dropout module. Variations in the probability of Dropout $p$ increased from 1×1$\times$1 × to 3×3$\times$3 × and the baseline model is the ideal scenario without variation.
Topologie Baseline Trained w/ Variations Trained w/o Variations
Var. 1×1$\times$1 × Var. 2×2$\times$2 × Var. 3×3$\times$3 × Var. 1×1$\times$1 × Var. 2×2$\times$2 × Var. 3×3$\times$3 ×
ResNet-18 91.52% 91.78% 91.77% 91.71% 91.65% 91.59% 91.58%
VGG 90.45% 90.52% 90.52% 90.55% 90.28% 90.26% 90.30%
In the other case, the NN is trained considering the variation in the Dropout module (see Section V-A). In this case, unlike in the previous case, there is a slight improvement in predictive performance (up to $+0.26\%$) for both models compared to the baseline. Variation in the Dropout probability leads to more stochastically during Bayesian inference, and as a result, accuracy improves slightly. The results are summarized in Table IV.
VII-B Uncertainty Estimation
Figure 5: Detecting Distribution Shift on CIFAR-10: a) A scatter and b) 95% confidence interval of 100 forward passes of the softmax input (logits) and output for Scale-Dropout VGG topology. Uniform noise of increasing strength is added to a randomly sampled image of a ship (leveled as 8). The uncertainty of the prediction increases with the data distribution shift as shown by the high SoftMax scatter and the confidence interval. Although the model uncertainty is extremely high (best observed in color), the input for images 5 through 12 is classified as either a truck (leveled as 9) or a bird (leveled as 3).
Figure 6: The effect of distribution shift of inference images on inference accuracy (left y-axis) and predictive entropy (uncertainty estimate on right y-axis). Images are continuously rotated to introduce distribution shifts. The inference accuracy of all methods reduces with the same trend and the uncertainty of prediction increases with the data distribution shift. The uncertainty estimates of the proposed method outperform those of other methods, but the accuracy of the Ensemble method is higher in comparison.
TABLE V: The analysis of the proposed Scale-Dropout BayBNN on three biomedical segmentation datasets on SOTA topologies. A Dropout probability of $20\%$ is used with $20$ Monte Carlo samples for Bayesian inference. The best-performing matrices are in bold.
U-Net Skin Cancer FP 32/32 95.10% 82.55% 98.81% 90.44% 89.04% 97.23% 91.88%
MC-Dropout [10] 32/32 95.05% 81.67% 98.86% 89.91% 84.74% 98.67% 95.74%
Proposed 1/4 96.75% 88.07% 99.4% 93.66% 92.31% 98.31% 95.04%
Bayesian SegNet Breast Cancer FP 32/32 97.47% 67.65% 96.60% 80.71% 76.95% 98.99% 84.84%
MC-Dropout [10] 32/32 97.83% 71.82% 96.40% 83.21% 78.23% 99.28% 88.86%
Spatial-SpinDrop [16] 1/4 96.32% 56.12% 94.31% 71.90% 68.50% 98.37% 75.64%
SpinDrop [17, 15] 1/4 97.47% 68.17% 95.48% 80.13% 74.25% 99.18% 87.02%
Proposed 1/4 97.69% 69.69% 96.69% 82.14% 77.19% 99.21% 87.76%
Bayesian SegNet COVID-19 Lung CT FP 32/32 99.53% 72.86% 99.11% 84.30% 83.17% 99.78% 85.46%
MC-Dropout [10] 32/32 99.51% 72.88% 99.39% 84.32% 86.99% 99.7% 81.8%
Proposed 1/4 99.55% 74.43% 99.69% 85.34% 85.84% 99.76% 84.85%
Figure 8: The outcomes of semantic segmentation and uncertainty estimations for the Skin-Cancer, COVID-19 Lung-CT, and Breast Cancer datasets. Each row comprises the input image followed by the ground truth, prediction mask, correctness mask, and uncertainty mask for both the MC-Dropout and proposed MC-Scale Dropout methods. The correctness mask is a binary representation of accurate and inaccurate predictions. The uncertainty mask is the normalized [00, $1$] uncertainty mask derived from twenty Monte Carlo samples. For prediction masks, a $0.5$ threshold is used. On the correctness and uncertainty masks, the correct and certain regions are depicted in white. Similarly, a region that is incorrect or uncertain is displayed in black.
Detecting Distribution Shift
On the other hand, we have performed another experiment with all images of the CIFAR-10 dataset on the VGG model continuously rotated up to $90^{\circ}$. It can be seen in Fig. 6 that as the images are rotated, the inference accuracy decreases and the predictive entropy increases from the starting entropy. Our method is compared with deterministic as well as common uncertainty estimation techniques, namely MC-Dropout [10] and Deep Ensemble [8] with five randomly initialized models. The trend of decrease in inference accuracy is similar for all models. However, the Deep Ensemble slightly outperforms the proposed and other methods in terms of accuracy. Regardless, our proposed MC Scale-Dropout method produces significantly more predictive entropy compared to other methods, including the Deep Ensemble method. This is because the proposed MC Scale-Dropout method effectively turns a single model into numerous ensembles by enabling Dropout during inference, allowing it to generate multiple predictions from a single model. Whereas, Deep Ensemble has limited models in the ensemble, e.g., five models. Also, the scale Dropout provides a regularization effect that can sometimes lead to better generalization in MC Scale-Dropout models compared to individual models in a Deep Ensemble. Consequently, our method can produce better uncertainty estimates compared to related works even with a 1-bit model.
Note that in the MC-Dropout model, Dropout is applied to the extracted features from the convolutional layers to achieve similar inference accuracy. If Dropout is applied to all layers, the predictive entropy increases, but inference accuracy decreases significantly, e.g., by more than 3%. Our approach provides a good balance between uncertainty estimates without any degradation in accuracy.
TABLE VI: Evaluation of the proposed MC-Scale Dropout method in detecting OOD across various topologies. All the models are trained on CIFAR-10 dataset.
VGG Proposed ${95.36\%}$ ${95.40\%}$ $\mathbf{96.19\%}$ $\mathbf{88.73\%}$ $\mathbf{98.02\%}$ ${85.24\%}$
Spatial-SpinDrop [16] $\mathbf{99.99\%}$ $\mathbf{100\%}$ $92.9\%$ $78.91\%$ $99.81\%$ $\mathbf{100\%}$
SpinBayes [14] $99.86\%$- $94.35\%$- $80.33\%$ $89.31\%$
ResNet-18 Proposed $\mathbf{100\%}$ $\mathbf{100\%}$ ${97.49\%}$ ${77.77\%}$ ${91.53\%}$ ${78.61\%}$
Spatial-SpinDrop [16] $100\%$ $100\%$ $\mathbf{100\%}$ $\mathbf{92.26\%}$ $\mathbf{99.98\%}$ $\mathbf{97.39\%}$
ResNet-20 Proposed ${96.51\%}$ $\mathbf{100\%}$ ${90.34\%}$ $\mathbf{93.55\%}$ $\mathbf{100\%}$ ${99.8\%}$
Spatial-SpinDrop [16] $\mathbf{97.2\%}$ $100\%$ $\mathbf{90.79\%}$ $87.94\%$ $99.03\%$ $\mathbf{99.81\%}$
Detecting Out-of-distribution Data
As depicted in Table VI, our proposed method can achieve a detection rate of OOD of up to $100\%$ across various model architectures and six different OOD datasets ($\hat{\mathcal{D}}_{1}$ through $\hat{\mathcal{D}}_{6}$). There are some variations in OOD detection rates across different architectures for the same OOD dataset. However, even in these cases, our method can consistently achieve a high OOD detection rate, with the lowest detection rate being $77.77\%$ on the ResNet-18 model with $\hat{\mathcal{D}}_{4}$ dataset. However, when the threshold for SoftMax confidence increases from $95\%$ to $99\%$, the OOD detection rate in the dataset $\hat{\mathcal{D}}_{4}$ improved to $81.78\%$, an $\sim 4\%$ improvement. Compared to MC-Spatial Dropout and SpinBayes methods, the OOD detection rates are generally similar. In the worst case, the OOD detection rate is $\sim 14\%$ lower for the VGG topology on the $\hat{\mathcal{D}}_{6}$ dataset. Therefore, the results indicate that the proposed MC-Scale Dropout method is a robust and reliable solution to OOD detection across diverse model architectures and datasets.
Epistemic Uncertainty of Semantic Segmentaion
VII-C Hardware Overhead Analysis